OpenStudy (itsbribro):

classify each scale factor as a contraction or expansion. 5 1/6 7/2 1/10 8/5 3/8

OpenStudy (itsbribro):

@LunyMoony

OpenStudy (lunymoony):

5 = 25 1/6 = 6/36 7/2 = 14/4 1/10 = 10/100 8/5 = 40/25 3/8 = 24/64 That's what my understand it anyway. Is that what you wanted or no? Or, you wanted a graph? There really was only a possiblity of expansion and not contraction to this

OpenStudy (mathstudent55):

Since when is 5 = 25?

OpenStudy (mathstudent55):

If the number is greater than 1, it is an expansion. If the number is less than 1, it is a contraction. 5 > 1 1/6 < 1 What about the others? Are they greater than or less than 1?

OpenStudy (itsbribro):

7/2> 1

OpenStudy (itsbribro):

@mathstudent55

OpenStudy (mathstudent55):

Correct, so 7/2 is an expansion.

OpenStudy (itsbribro):

1/10 is a contraction 3/8 is a contraction 8/5 is an expansion am i right?

OpenStudy (mathstudent55):

A positive whole number other than 1 is always greater than 1. Dealing with positive numerators and denominators: A fraction with a numerator greater than the denominator is an improper fraction and is also greater than 1. A fraction with a numerator less than the denominator is less than 1.

OpenStudy (mathstudent55):

You are correct.

OpenStudy (itsbribro):

yayy thank you so much